Order: NS 500ml to be infused over 12 hours a. what rate will infusion pump be set at b. how much NS will infuse after 4.5 hours
Answer #1

Ans) Order: NS 500ml to be infused

Time: 12 hours

a. Infusion rate (ml/hr) : Total volume to be infused ÷ time in hr

= 500 ÷ 12 = 41.66 ml/hr

The Infusion rate is to be set at 41.66 ml/hr.

b. after 4.5 hrs,

41.66 × 4.5 = 187.5 ml.

18.75 ml will be infused after 4.5 hrs.